The sum of a 2 digit number is 8. When 18 is added to that number the digits get interchanged. Find the number. (7th Std-CBSE)
Answer by MRperkins(300)

The sum of a 2 digit number is 8:
possibilities=
1+7 or 17
2+6 or 26
3+5 or 35
4+4 or 44
5+3 or 53
6+2 or 62
7+1 or 71
When 18 is added to that number the digits get interchanged. Find the number.
17+18=35 False
26+18=44 False
35+18=53 True
44+18=62 False
53+18=71 False
62+18=80 False
71+18=89 False
This is solved using a truth table.